OpenStudy (anonymous):

Come on Calculus Friends!! If you are given curl i.e. (del (cross) F) what is the method to figure out what F is, when (del (cross) F) is a vector quantity i.e. (#, #, #)

OpenStudy (anonymous):

there isn't an easy way to do it

OpenStudy (anonymous):

you've got yourself a lovely set of 3 partial differential equations. \[\frac{\delta u_y}{\delta z}-\frac{\delta u_z}{\delta y}=f(x,y,z) \hat{x}\] etc.

OpenStudy (anonymous):

can I show you the actual question and get your opinion on what to do from there?

OpenStudy (anonymous):

yeah, that'd be helpful, cause in general there's not really a set method of doing it that I'm aware of

OpenStudy (anonymous):

Let F and G be vector fields such that (del) X F(0) = (-6.914, 5.159, 4.502), G(0) = (-8.196, -1.838, -2.2). Find the divergence of F X G at 0.

OpenStudy (anonymous):

as far as i know the identity is shown as : (del)∗(FXG)=((del)crossF)dotG)−Fdot((del)crossG)

OpenStudy (anonymous):

ah, okay, this is a bit of a trick, the divergence of a curl is always zero.

OpenStudy (anonymous):

Ha! problem solved

OpenStudy (anonymous):

oh wait, nvm, that wasnt in the question...

OpenStudy (anonymous):

well the second part it kind of is no?

OpenStudy (anonymous):

no, its the dot product with a curl, its different.

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

so you are saying that F(dot)[del(cross)G) is not.. a divergence of a curl?

OpenStudy (anonymous):

I guess, yeah that wouldn't make much sense, because then why would they have that identity in the first place

OpenStudy (anonymous):

.. if it was just zero

OpenStudy (anonymous):

yeah, g(0) is all constants though, shouldnt the curl of it be zero?

OpenStudy (anonymous):

so all you'll have is the dot product of the two? I could be wrong, but from what I see you're definitly not given enough information to reverse engineer the field for f out of what you have

OpenStudy (anonymous):

See, those are the types of things i'm still trying to grasp.. aha. it would be nice make things much easier

OpenStudy (anonymous):

okay then that must be the best bet, I have possible answers, so i'll try and if I get it I'll let you know either way. thank you for your help!

OpenStudy (anonymous):

k, best of luck